Fermi National Laboratory
 

Accelerator Update for October 17, 2007 - October 18, 2007

 

Wednesday October 17

The day shift began with the accelerator complex in startup. SciBooNE / MiniBooNE took beam when available. The Main Injector (MI), Antiproton Source (Pbar), and Tevatron (TeV) all conducted studies.

Crew Chief Summary: The day shift started with a Linac RF station (LRF2) problem that wasn’t resolved until 2 PM. During this downtime, experts accessed the TeV sectors A-E and the Transfer Hall to adjust lead flows. When beam returned, Linac experts started working on LRF5 and a kicker power supply (KPS4A), which caused spotty beam.

The evening shift began with studies on hold due to the KPS4A problems. Beam resumed at 5:30 PM. Pbar experts worked at phasing in cooling until 11:30 PM. They reported that they would need to access the rings to repair the Pbar Debuncher momentum cooling. Experts worked on tuning the MI8 transfer line. Recycler experts completed their studies and turned off the Pelletron for the night. MI (MI-30) suffered a vacuum burst; it recovered. The TeV expert studies focused on 150 GeV beam quality and on reverse proton injection. They reported some problems with their controls system, on a TeV front-end. Controls experts worked on the ACNET alarms problems.

Thursday October 18

The midnight shift began with the accelerator complex in startup, with TeV experts working on a separator problem, with the Recycler conducting studies, and with Sci/MiniBooNE taking beam.

Crew Chief Summary: A Recycler expert conducted transfer line studies until 6 AM. Operations tuned the A150 transfer line in order to get reverse protons circulating in MI. TeV experts began conducting studies by 4 AM. A quadrupole power supply in the P150 transfer line tripped off regularly; operators believed this was due to a readback problem

The Plans for Thursday

The plans for today are to allow Pbar a four-hour access to their rings and work on their Debuncher momentum cooling. All machines will continue their commissioning efforts.

Machine Summaries

Linac: Experts tuned on the nine-inch transmission line for LRF5 on Wednesday. They want to revert back to its previous state, which would take about one-hour of no beam. This will be done later.

Booster: Delivered beam to accelerators and the Sci/MiniBooNE (S/MB) experiments.

Main Injector: Experts smoothed the MI-8 transfer line, but think they might have caused beam problems for S/MB. They reported that they have DCCT problems and a problem with their closure program. The MI-8 line needs more work and they haven’t started working on coalescing.

Antiproton Source: Experts phased their stacktail and Debuncher systems. They conducted a target scan. After their access they want to conduct reverse proton studies, stack, and measure their tunes. They would like to stack all night tonight and then transfer antiprotons to the Recycler tomorrow.

Recycler: Experts conducted Electron Cooling and proton studies as well as sending protons back and forth between the Recycler and MI.

Tevatron: Experts reported that lead temperatures improved after they made their adjustments. They reported that they were having problems with their closure program. They would like to accelerate beam later this afternoon. They are not ready for coalesced beam or for antiprotons.

Cryogenic Systems: Everything was fine and cold.

Controls: Experts continue to work on the alarms problem.

D0: They worked on some small jobs during their access. They still want an 8-10 hour access. They were taking cosmic rays.

CDF: They worked on small jobs during their access. They found and fixed a ground line problem. They were taking cosmic rays.

FESS: An expert reported that Casey’s pond water level looked good. At CUB they found a cable fault on a Pbar water pump. As a temporary measure they will splice around the fault. This should fix it until they have an opportunity to pull a new cable. They completed their TeV sector D4 motor repair. They successfully conducted a fire pump test for Wilson Hall.

MiniBooNE/SciBooNE: They were taking beam and tuning.

SwitchYard 120: Shutdown continued.

NuMI: Shutdown continued.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/18/2007    email Fermilab